Write the vector (1, −2, 5) as a linear combination of the vectors (1, 1, 1),(1, 2, 3)

and (2, −1, 1)


1
Expert's answer
2022-02-14T10:31:32-0500


Let us write the vector "(1, \u22122, 5)" as a linear combination of the vectors "(1, 1, 1),(1, 2, 3)"

and "(2, \u22121, 1)."


Let

 "(1, \u22122, 5)=a(1, 1, 1)+b(1, 2, 3)+c(2, \u22121, 1)=(a+b+2c,a+2b-c,a+3b+c)."


It follows that we get the system


"\begin{cases}
a+b+2c=1\\
a+2b-c=-2\\
a+3b+c=5
\end{cases}"


which is equivalent after subtracting from the second row the first row, and from the third row the second row to the system


"\begin{cases}
a+b+2c=1\\
b-3c=-3\\
b+2c=7
\end{cases}"


and hence the last system is equivalent after subtracting from the third row the second row to the system


"\begin{cases}
a+b+2c=1\\
b-3c=-3\\
5c=10
\end{cases}"


It follows that

 "c=2,\\ b=3c-3=6-3=3,\\ a=1-b-2c=1-3-4=-6."


We conclude that "(1, \u22122, 5)=-6(1, 1, 1)+3(1, 2, 3)+2(2, \u22121, 1)."